OpenStudy (anonymous):

Container A has twice the capacity of container B. Container A is full of sand and container B is empty. suppos 1/8 of the sand in container A is poured into container B. What fractional part of container B will contain sand?

OpenStudy (anonymous):

A = 2B (1/8)A = (1/8)(2B) = (1/4)B (1/4)B/B = 1/4 <--